OpenStudy (anonymous):

Factoring and using the quadratic formula: 2y^2-3y-6=0

OpenStudy (marihelenh):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

using quadratic formula y=1+/-(1+4*2*6)^1/2 whole divided by 2.2 y=-3/2 and 2 y=

OpenStudy (anonymous):

using factorization 2y^2-y-6=0 2y^2-4y+3y-6=0 2y(y-2)+3(y-2)=0 (2y+3)(y-2)=0 answer=-3/2&2

OpenStudy (anonymous):

This is my best answer (2y+3)(y-2)

OpenStudy (anonymous):

2y² - y - 6 = 0 y = [1±√ (1 + 4 x 2 x 6)] / 4 y = [ 1± √49 ] / 4 y = 8/4 or -6/4 y = 2 or - 3/2 Can be checked by factorising:- 2y² - y - 6 = 0 (2y + 3).(y - 2) = 0 y = - 3/2 or y = 2 --------------(as above)